# Competition between transient oscillations and early stochasticity in exponentially growing populations

### v2

@@ I Introduction
The relationship between bacterial populations and single-cell statistics has extensively been researched, particularly for cells in unrestricted growth [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In the presence of sufficient nutrients, bacterial populations grow exponentially at a constant rate that can be linked to the single-cell statistics [3, 4, 5, 6, 7, 9, 14, 15, 12, 19, 24]. However, when the population starts from a few or a single cell, it takes some time before bacterial cell count grows perfectly exponentially, even if all individual cells are in balanced growth, meaning that their physiological states are adapted to the growth medium. This transient period is characterized by slowly decaying oscillations [20, 10, 11, 18, 13, 26, 1, 25]. These oscillations occur because the descendants of a cell grow together and divide more or less at the same time resulting in synchronized bursts of division in the population. As time passes, the stochasticity in cell growth slowly de-synchronizes the timing of cell division, which dampens the oscillations [1, 25]. Note that these oscillations only show in population observables that are affected by divisions, such as total cell count, cell count growth rate, or cell mass distribution. The total cell mass grows smoothly, which is why optical density measurements would not reveal the oscillations [25].
@@
To illustrate how transient behavior manifests, we focus on the cell count growth rate $\Lambda(t)$, defined as the fraction of cells in the population that divide per unit time $\frac{1}{N(t)}\frac{dN(t)}{dt}$. This population observable, along with many others, converges to a certain asymptotic constant value after an infinite amount of time. At large but finite times, however, these population observables exhibit transient behavior. The ensemble averages of intensive observables oscillate around their asymptotic value, with an exponentially decaying amplitude. The time scale for this oscillation decay has been derived for several single-cell models of growth and division, and it scales with the variability of single-cell growth rates [20, 10, 11, 18, 13, 26, 1, 25]. On the other hand, due to stochasticity in single-cell growth and division, cell count $N(t)$ evolves stochastically with fluctuations that scale with $\sqrt{N(t)}$. The relative amplitude of such fluctuations, therefore, scales as $1/\sqrt{N(t)}$, which decays exponentially at a rate of half the population growth rate. In other words, both the oscillations and fluctuations in cell count growth rate decay exponentially with competing decay rates.
@@
There is a transition point at which the transient oscillations and fluctuations are equal. Since the oscillations and fluctuations decrease in magnitude at different rates, the transient population behavior will ultimately be dominated by only one of these effects: the one with the lowest decay rate. We can, therefore, formulate a condition necessary for the oscillations to dominate the transient population behavior. In Ref. [1], the oscillation decay rate was estimated to be $r\approx 29\,\sigma_{\kappa}^{2}/\bar{\kappa}$, where $\sigma_{\kappa}$ and $\bar{\kappa}$ are the single-cell growth rate standard deviation and mean respectively. The asymptotic population growth rate is roughly identical to the single-cell growth rate $\bar{\kappa}$, so the fluctuations decay at a rate of $\bar{\kappa}/2$. We can therefore say that oscillations dominate the transient behavior only if
@@
This transition point lies within the range of physiological values for $\sigma_{\kappa}$ and $\bar{\kappa}$ [27].
The aforementioned transition point was defined for a model in which growth rate correlations were neglected. Population behavior depends on how cell division times are distributed and how they are correlated between mother and daughter cells [5, 7]. A good way to model these correlations is by treating growth rate variability and division size variability as two separate sources of noise, each with its own mother-daughter cell correlations Ref. [27]. In Ref. [25] the authors use such a description of single-cell behavior to quantitatively predict the oscillatory population behavior. They assume that cell growth rate along a lineage is a continuous random process. The per-cell growth rate variability and correlations then emerge from the noise and auto-correlation of the underlying growth process. All of their results have, however, only been derived for ensemble averaged populations, which means they only considered observables corresponding to averages over repeated experiments with the same initial conditions.
In this paper, we adopt the model from Ref. [25], but consider the behavior of single population experiments instead of just the ensemble average, which causes some interesting effects that are often ignored in other literature. Unlike in the ensemble average population, we observe a transition between fluctuation-dominated and oscillation-dominated transient regimes. We determine a generalized transition point, of which Eq. (1) is a special case. Another interesting phenomenon absent in the ensemble average is that the amplitude of the transient oscillations is random among repeated experiments, set by stochasticity in the early population. We analyze the distribution of these transient oscillation amplitudes through simulations and theory. Interestingly, the typical amplitude of a single population is higher than that of an ensemble-averaged population, due to the net effects of spontaneous cell synchronization. Our predictions are in great qualitative agreement with oscillations studied in a multistage model used to describe cancer cell populations [11]. At last, we consider tables of realistic parameters based on data from Ref. [27] and answer just how many generations it typically takes for the oscillations to fall below a certain threshold.
@@ II Model
Suppose we have a population that starts from a single cell. We assume this cell is already adapted to the growth medium. We want to keep track of the number of cells $N(t)$. For this, we use a physiologically relevant model of growth and division, which accounts for growth fluctuations, their correlations, and cell size regulation based on the model from Ref. [25]. The effect of asymmetric division is ignored, since its effect of oscillations is relatively small for physiologically relevant values [1, 28, 29, 30].
@@
$\frac{d}{dt}p(t,\mathbf{x})=\mathcal{K}p(t,\mathbf{x}),$ (3)
@@
In our simulations, we assume that cell masses at division $m_{d}$ depend on their masses at birth $m_{b}$, via the relation based on the cell size regulation model from Ref. [31],
@@
where $\bar{m}$ is the equilibrium cell birth mass, $\alpha$ is the mother-daughter division mass correlation parameter and $\eta$ is a normally distributed random variable with mean zero and variance $\sigma_{\eta}^{2}$. We assume symmetric division, so the daughter cells’ birth masses $m_{d}$ are set to be half their mother cell’s mass at the division. After a few generations, the mass at birth and division quickly settle to equilibrium log-normal distributions with $\langle\ln m_{b}\rangle=\ln\bar{m}$ and $\langle\ln m_{d}\rangle=\ln 2\bar{m}$ respectively [25]. Their coefficients of variation (ratio of standard deviation to mean) are identical and given by
@@
In our simulations, we always use $\alpha=0.5$, which corresponds to an adder model of cell size regulation [31, 27].
We would like to emphasize that our analytical results do not assume this particular choice of cell size regulation model used in the simulation. We will show that the only aspect of the cell division model that affects long-term population dynamics is the value of $\text{CV}_{m_{d}}$.
@@ III Ensemble averaged oscillations and Oscillation-fluctuation trade-off
This effective growth rate is known to converge to an asymptotic population growth rate $\Lambda_{\infty}$ at large times [9, 5, 7]. This limit holds true when the cell mass and growth state distribution of the population have reached a steady state, in which case the cell masses have completely de-synchronized.
Let us explain why a non-steady-state population exhibits oscillations in population growth rate. When the average cell mass exceeds the steady-state average, we have an increase in cells close to division, which momentarily raises $\Lambda(t)$ with respect to the steady-state value $\Lambda_{\infty}$. Following an increase in divisions, the abundance of newborn cells causes average cell mass to drop below the steady state, thereby decreasing $\Lambda_{\text{ens}}(t)$. When the number of divisions $\Lambda_{\text{ens}}(t)$ is low, average cell mass builds up again and the cycle repeats. This process causes wave-like behavior in cell mass distribution and cell division rate $\Lambda_{\text{ens}}(t)$. The average effect of cells growing at different rates de-synchronizes the population. This dampens the mass synchronicity waves, at rates proportional to their oscillation speed. This effect rapidly dampens out any higher-order mass synchronicity oscillations. What is left are just the first-order oscillations corresponding to full cell cycles. The transient regime is thus well described by just the first-order oscillations  [25]
@@
Here, the amplitude pre-factor $A_{\text{ens}}$ and phase $\phi_{\text{ens}}$ are constants that depend on the division process, growth process as well as the population’s initial conditions. An explicit expression for $A_{\text{ens}}$ is given in Section IV. The asymptotic growth rate $\Lambda_{\infty}$, oscillation decay rate $r$, and oscillation speed $\Omega$, however, are time scales that are uniquely determined by the model of cell growth. In Ref. [25] these timescales were given in the case where $\lambda_{t}$ is an Ornstein-Uhlenbeck process. Here, we provide a more general expression.
Assuming a growth rate process $\lambda(\mathbf{x})$ where $\mathbf{x}$ evolves according to Eq. (3), we find that $\Lambda_{\infty}$ is the leading eigenvalue of the operator $\mathcal{K}+\lambda(\mathbf{x})$. Let $\mu$ be the leading eigenvalue by real part of the operator $\mathcal{K}+[1+i2\pi/\ln(2)]\lambda(\mathbf{x})$. Then we have that $r=\Lambda_{\infty}-\text{Re}\mu$ and $\Omega=\text{Im}\mu$. These expressions reduce to forms that are easier to interpret if we assume $\lambda_{t}$ is Ornstein-Uhlenbeck or assume small growth rate variability $\sigma_{\lambda}\ll\bar{\lambda}$. For any growth rate process, one can define an auto-correlation function
@@
For a growth process $\lambda_{t}$ that is Ornstein-Uhlenbeck or has small growth rate variability $\sigma_{\lambda}\ll\bar{\lambda}$, the population time scales are given by are given by
  • $\Lambda_{\infty}=\bar{\lambda}+D$
  • $r=\frac{4\pi^{2}}{\ln(2)^{2}}D$
  • $\Omega=\frac{2\pi}{\ln(2)}(\bar{\lambda}+2D)$,
where $D$ is given by Eq. (9). When $\lambda_{t}$ is Ornstein-Uhlenbeck, we have that $D=\sigma_{\lambda}^{2}\tau_{\text{cor}}$.
Figure 1: Population growth rate as a function of time for two parameter values on both sides of the transition. (a) growth rate above the transition is stochastic and approaches steady exponential growth, while (c) the growth rate below the transition approaches a decaying oscillation around the exponential growth. (b) and (d) are the same plots in logscale (with asymptotic values subtracted, $\ln_{>0}[(\Lambda(t)-\Lambda_{\infty})/\bar{\lambda}]$) showing the exponential decay of transient dynamics. In (b), above the transition, oscillations decay faster than the fluctuations, while in (d) the oscillations decay slower than the fluctuations and become asymptotically deterministic. Simulation parameters: $\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}=0.008$ for (a) and (b), and $\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}=0.003$ for (c) and (d).
@@
Suppose that the $N(t)$ cells in a population grow independently. The fluctuations in the number of cells that divide per unit time $dN(t)/dt$ now scale with $\sqrt{N(t)}$. Consequently, the fluctuations in the cell count growth rate $\Lambda(t)$ scale with $1/\sqrt{N(t)}$, which evolves as $e^{-\frac{\Lambda_{\infty}}{2}t}$. This introduces a timescale of $\Lambda_{\infty}/2$ in addition to the oscillation decay rate $r$. We define the relative oscillation decay rate $\epsilon$ as the ratio of the oscillation decay rate to the fluctuation decay rate
$\epsilon:=\frac{r}{\Lambda_{\infty}/2}\approx 164D/\bar{\lambda}.$ (11)
If $\epsilon<1$, then the oscillations decay more slowly than the background fluctuations. In this regime, the ratio of fluctuations to oscillations in $\Lambda(t)$ will always go to zero eventually. When this happens, the dynamics of $\Lambda(t)$ become deterministic. Conversely, if $\epsilon\geq 1$, then all oscillations will eventually fall below the background fluctuation level. the trajectory of $\Lambda(t)$ is dominated by noise and local oscillations caused by spontaneous cell mass synchronicity. In Appendix D we show how in the limit of vanishing growth rate correlations, the condition $\epsilon<1$ is equivalent to Eq. (1).
In Fig. 1 we show examples of the trajectories of $\Lambda(t)$ for $\epsilon=0.5$ and $\epsilon=1.3$, as well as the log of the deviation of $\Lambda(t)$ from its asymptotic value $\Lambda_{\infty}$. In log deviation space, the exponential decay rates of transient effects appear as slopes. Which transient effect dominates in the large time limit only depends on the theoretical ratio of the slopes set by $\epsilon$. One can think of the point at which the oscillations clearly separate from the background noise as the crossover between the two lines. At the transition, as $\epsilon$ goes to $1$, this separation time diverges. In Sec. IV we will show that this dynamic transition at $\epsilon=1$ has properties that are analogous to equilibrium phase transitions.

### v3

@@ I Introduction
The relationship between bacterial populations and single-cell statistics has extensively been researched, particularly for cells in unrestricted growth Jafarpour (2019); Jafarpour et al. (2018); Hashimoto et al. (2016); Levien et al. (2020a); Lin and Amir (2017); Ho et al. (2018); Lin and Amir (2020); Luo et al. (2021); Powell (1956); Nozoe and Kussell (2020); Gavagnin et al. (2021); Thomas (2017); Chiorino et al. (2001); Tănase-Nicola and Ten Wolde (2008); Thomas (2018); García-García et al. (2019); Levien et al. (2020b); Vittadello et al. (2019); Marantan and Amir (2016); Kendall (1948); Hirsch and Engelberg (1966); Genthon (2022); Barber et al. (2021); Jia et al. (2021); Hein and Jafarpour (2022). In the presence of sufficient nutrients, bacterial populations grow exponentially at a constant rate that can be linked to the single-cell statistics Hashimoto et al. (2016); Levien et al. (2020a); Lin and Amir (2017); Ho et al. (2018); Lin and Amir (2020); Powell (1956); Tănase-Nicola and Ten Wolde (2008); Thomas (2018, 2017); Marantan and Amir (2016); Jia et al. (2021). However, when the population starts from a few or a single cell, it takes some time before bacterial cell count grows perfectly exponentially, even if all individual cells are in balanced growth, meaning that their physiological states are adapted to the growth medium. This transient period is characterized by slowly decaying oscillations Kendall (1948); Nozoe and Kussell (2020); Gavagnin et al. (2021); Vittadello et al. (2019); Chiorino et al. (2001); Yates et al. (2017); Jafarpour (2019); Hein and Jafarpour (2022). These oscillations are distinct from the type of oscillations that arise due to cell interactions, such as resource competitions and stress signals Martinez-Corral et al. (2018). Instead, these oscillations occur because the descendants of a cell grow together and divide more or less at the same time resulting in synchronized bursts of division in the population. As time passes, the stochasticity in cell growth slowly de-synchronizes the timing of cell division, which dampens the oscillations Jafarpour (2019); Hein and Jafarpour (2022).
@@
To illustrate how transient behavior manifests, we focus on the cell count growth rate $\Lambda(t)$, defined as the fraction of cells in the population that divide per unit time $\frac{1}{N(t)}\frac{dN(t)}{dt}$. This population observable, along with many others, converges to a certain asymptotic constant value after an infinite amount of time. At large but finite times, however, these population observables exhibit transient behavior. The ensemble averages of intensive observables oscillate around their asymptotic value, with an exponentially decaying amplitude. The time scale for this oscillation decay has been derived for several single-cell models of growth and division, and it scales with the variability of single-cell growth rates Kendall (1948); Nozoe and Kussell (2020); Gavagnin et al. (2021); Vittadello et al. (2019); Chiorino et al. (2001); Yates et al. (2017); Jafarpour (2019); Hein and Jafarpour (2022). On the other hand, due to stochasticity in single-cell growth and division, cell count $N(t)$ evolves stochastically with fluctuations that scale with $\sqrt{N(t)}$. The relative amplitude of such fluctuations, therefore, scales as $1/\sqrt{N(t)}$, which decays exponentially at a rate of half the population growth rate. In other words, both the oscillations and fluctuations in cell count growth rate decay exponentially with competing decay rates.
@@
There is a transition point at which the transient oscillations and fluctuations are equal. Since the oscillations and fluctuations decrease in magnitude at different rates, the transient population behavior will ultimately be dominated by only one of these effects: the one with the lowest decay rate. We can, therefore, formulate a condition necessary for the oscillations to dominate the transient population behavior. In Ref. Jafarpour (2019), the oscillation decay rate was estimated to be $r\approx 29\,\sigma_{\kappa}^{2}/\bar{\kappa}$, where $\sigma_{\kappa}$ and $\bar{\kappa}$ are the single-cell growth rate standard deviation and mean respectively. The asymptotic population growth rate is roughly identical to the single-cell growth rate $\bar{\kappa}$, so the fluctuations decay at a rate of $\bar{\kappa}/2$. We can therefore say that oscillations dominate the transient behavior only if
@@
This transition point lies within the range of physiological values for $\sigma_{\kappa}$ and $\bar{\kappa}$ Taheri-Araghi et al. (2015).
The aforementioned transition point was defined for a model in which growth rate correlations were neglected. Population behavior depends on how cell division times are distributed and how they are correlated between mother and daughter cells Lin and Amir (2017, 2020). A good way to model these correlations is by treating growth rate variability and division size variability as two separate sources of noise, each with its own mother-daughter cell correlations Taheri-Araghi et al. (2015). In Ref. Hein and Jafarpour (2022) the authors use such a description of single-cell behavior to quantitatively predict the oscillatory population behavior. They assume that cell growth rate along a lineage is a continuous random process. The per-cell growth rate variability and correlations then emerge from the noise and auto-correlation of the underlying growth process. All of their results have, however, only been derived for ensemble averaged populations, which means they only considered observables corresponding to averages over repeated experiments with the same initial conditions.
In this paper, we adopt the model from Ref. Hein and Jafarpour (2022), but consider the behavior of single population experiments instead of just the ensemble average, which causes some interesting effects that are often ignored in other literature. Unlike in the ensemble average population, we observe a transition between fluctuation-dominated and oscillation-dominated transient regimes. We determine a generalized transition point, of which Eq. (1) is a special case. Another interesting phenomenon absent in the ensemble average is that the amplitude of the transient oscillations is random among repeated experiments, set by stochasticity in the early population. We analyze the distribution of these transient oscillation amplitudes through simulations and theory. Interestingly, the typical amplitude of a single population is higher than that of an ensemble-averaged population, due to the net effects of spontaneous cell synchronization. Our predictions are in great qualitative agreement with oscillations studied in a multistage model used to describe cancer cell populations Gavagnin et al. (2021). At last, we consider tables of realistic parameters based on data from Ref. Taheri-Araghi et al. (2015) and answer just how many generations it typically takes for the oscillations to fall below a certain threshold.
@@ II Model
Suppose we have a population that starts from a single cell. We assume this cell has already been adapted to the growth medium. We want to keep track of the number of cells $N(t)$. For this, we use a physiologically relevant model of growth and division, which accounts for growth fluctuations, their correlations, and cell size regulation based on the model from Ref. Hein and Jafarpour (2022). The effect of asymmetric division is ignored, since its effect of oscillations is relatively small for physiologically relevant values Jafarpour (2019); Eun et al. (2018); Campos et al. (2014); Guberman et al. (2008).
@@
$\frac{\partial}{\partial t}p(t,\mathbf{x})=\mathcal{K}p(t,\mathbf{x}),$ (3)
@@
When concrete models of division are needed, we assume that cell masses at division $m_{d}$ depend on their masses at birth $m_{b}$, via the relation based on the cell size regulation model from Ref. Amir (2014),
@@
where $\alpha$ is the mother-daughter division mass correlation parameter, $\eta$ is a normally distributed random variable with mean zero and variance $\sigma_{\eta}^{2}$, and $\bar{m}$ is a constant cell mass that corresponds to an equilibrium birth mass. We assume symmetric division, so the daughter cells’ birth masses $m_{d}$ are set to be half their mother cell’s mass at the division. After a few generations, the mass at birth and division quickly settle to equilibrium log-normal distributions with $\langle\ln m_{b}\rangle=\ln\bar{m}$ and $\langle\ln m_{d}\rangle=\ln 2\bar{m}$ respectively Hein and Jafarpour (2022). This motivates calling $\bar{m}$ the equilibrium birth mass. The coefficients of variation of birth mass and division mass are identical and given by
@@
In our simulations, we always use $\alpha=0.5$, which corresponds to an adder model of cell size regulation Amir (2014); Taheri-Araghi et al. (2015). The simulation code is available at Cod
We emphasize that our analytical results do not assume this particular choice of cell size regulation model used in the simulation. We will show that the only aspect of the cell division model that affects long-term population dynamics is the value of $\text{CV}_{m_{d}}$.
@@ III Ensemble averaged oscillations and Oscillation-fluctuation trade-off
This effective growth rate is known to converge to an asymptotic population growth rate $\Lambda_{\infty}$ at large times Powell (1956); Lin and Amir (2017, 2020). This limit holds true when the cell mass and growth state distribution of the population have reached a steady state, in which case the cell masses have completely de-synchronized. Alternatively, we could define the population growth rate as the change in the average logarithm of population count $d\left\langle\ln N(t)\right\rangle/dt$ Lewontin and Cohen (1969). We use Eq. (6) because it is easier to work with analytically while it still exhibits the same oscillatory transient behavior of interest.
Let us explain why a non-steady-state population exhibits oscillations in population growth rate. When the average cell mass exceeds the steady-state average, we have an increase in cells close to division, which momentarily raises $\Lambda(t)$ with respect to the steady-state value $\Lambda_{\infty}$. Following an increase in divisions, the abundance of newborn cells causes average cell mass to drop below the steady state, thereby decreasing $\Lambda_{\text{ens}}(t)$. When the number of divisions $\Lambda_{\text{ens}}(t)$ is low, average cell mass builds up again and the cycle repeats. This process causes wave-like behavior in cell mass distribution and cell division rate $\Lambda_{\text{ens}}(t)$. The average effect of cells growing at different rates de-synchronizes the population. This dampens the mass synchronicity waves, at rates proportional to their oscillation speed. This effect rapidly dampens out any higher-order mass synchronicity oscillations. What is left are just the first-order oscillations corresponding to full cell cycles. The transient regime is thus well described by just the first-order oscillations  Hein and Jafarpour (2022)
@@
Here, the amplitude pre-factor $A_{\text{ens}}$ and phase $\phi_{\text{ens}}$ are constants that depend on the division process, growth process as well as the population’s initial conditions. An explicit expression for $A_{\text{ens}}$ is given in Section IV. The asymptotic growth rate $\Lambda_{\infty}$, oscillation decay rate $r$, and oscillation speed $\Omega$, however, are time scales that are uniquely determined by the model of cell growth. In Ref. Hein and Jafarpour (2022) these timescales were given in the case where $\lambda_{t}$ is an Ornstein-Uhlenbeck process. Here, we provide a more general expression.
Assuming a growth rate process $\lambda(\mathbf{x})$ where $\mathbf{x}$ evolves according to Eq. (3), we find that $\Lambda_{\infty}$ is the leading eigenvalue of the operator $\mathcal{K}+\lambda(\mathbf{x})$. Let $\mu$ be the leading eigenvalue by real part of the operator $\mathcal{K}+[1+i2\pi/\ln(2)]\lambda(\mathbf{x})$. Then we have that the oscillation decay rate is $r=\Lambda_{\infty}-\text{Re}\mu$ and the oscillation speed is $\Omega=\text{Im}\mu$. These expressions reduce to forms that are easier to interpret if we assume $\lambda_{t}$ is Ornstein-Uhlenbeck or assume small growth rate variability $\sigma_{\lambda}\ll\bar{\lambda}$. For any growth rate process, one can define an auto-correlation function
@@
This parameter acts as the effective diffusion constant of the integrated growth rate process over time scales much larger than the correlation time $\tau_{\text{cor}}$. For a growth process $\lambda_{t}$ that is Ornstein-Uhlenbeck or has small growth rate variability $\sigma_{\lambda}\ll\bar{\lambda}$, the population time scales are
  • population growth rate $\Lambda_{\infty}=\bar{\lambda}+D$
  • oscillation decay rate $r=\frac{4\pi^{2}}{\ln(2)^{2}}D\approx 82D$
  • oscillation speed $\Omega=\frac{2\pi}{\ln(2)}(\bar{\lambda}+2D)$.
When $\lambda_{t}$ is Ornstein-Uhlenbeck, we have that $D=\sigma_{\lambda}^{2}\tau_{\text{cor}}$. A derivation of the time scales is shown in Appendix B. When considering collective oscillations, small phase shifts among the oscillations of sub-populations can greatly affect the amplitude of their sum due to destructive interference. This is reflected in the large constant $4\pi^{2}/\ln(2)^{2}\approx 82$ that connects the oscillation amplitude decay rate $r$ to growth fluctuations. Any large constants encountered in the rest of the text result from similar effects and their exact form can be found in the appendix.
Figure 1: Population growth rate as a function of time for two parameter values on both sides of the transition. (a) growth rate above the transition is stochastic and approaches steady exponential growth, while (c) the growth rate below the transition approaches a decaying oscillation around the exponential growth. (b) and (d) are the same plots in logscale (with asymptotic values subtracted, $\ln_{>0}[(\Lambda(t)-\Lambda_{\infty})/\bar{\lambda}]$) showing the exponential decay of transient dynamics. In (b), above the transition, oscillations decay faster than the fluctuations, while in (d) the oscillations decay slower than the fluctuations and become asymptotically deterministic. Simulation parameters: $\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}=0.008$, which is $\epsilon=1.3$ for (a) and (b), and $\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}=0.003$, which is $\epsilon=0.5$ for (c) and (d). All simulations start from a single cell with cell mass $m_{0}=\bar{m}=1$ and growth rate $\lambda_{0}=\bar{\lambda}=\ln(2)$ (such that $\tau_{\text{div}}=1$). Simulation details: Given a cell’s birth size, its target division size is randomized according to Eq. (4). with division parameters $\text{CV}_{m_{d}}=0.12$ and $\alpha=0.5$. At every time increment of $dt=0.01$, each cell’s growth rate growth rate is updated as an Ornstein-Uhlenbeck process, $\lambda_{t+dt}=e^{-dt/\tau_{\text{cor}}}\lambda_{t}+\sqrt{2\sigma_{\lambda}^{2}/\tau_{\text{cor}}}dW_{t}$, where $dW_{t}$ is drawn from a normal distribution such that $\mathbb{E}[dW_{t}]=0$ and $\text{Var}(dW_{t})=dt$. Cell masses are updated as $m_{t+dt}=m_{t}e^{\lambda_{t}dt}$. After every time step, for every cell whose mass now exceeds the division mass $m_{t}>m_{d}$, we take that cell out of the population and replace it by two daughter cells with birth size $m_{t}/2$ and the same growth rate $\lambda_{t}$ as their mother cell. Throughout the simulation, we track the number of cells $N(t)$ at every time step. Population growth rate $\Lambda_{t}$ is determined using Eq. (33) with a bin size of $\Delta t=0.1$. In Appendices C.1 and C.2 we discuss how we obtained the fluctuation and oscillation fit envelopes respectively. The simulation code is provided in the GitHub repository Cod .
@@
Suppose that the $N(t)$ cells in a population grow independently. The fluctuations in the number of cells that divide per unit time $dN(t)/dt$ now scale with $\sqrt{N(t)}$. Consequently, the fluctuations in the cell count growth rate $\Lambda(t)$ scale with $1/\sqrt{N(t)}$, which evolves as $e^{-\frac{\Lambda_{\infty}}{2}t}$. This introduces a timescale of $\Lambda_{\infty}/2$ in addition to the oscillation decay rate $r$. We define the decay ratio $\epsilon$ as the ratio of the oscillation decay rate to the fluctuation decay rate
$\epsilon:=\frac{r}{\Lambda_{\infty}/2}\approx 164\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}.$ (11)
If $\epsilon<1$, then the oscillations decay more slowly than the background fluctuations. In this regime, the ratio of fluctuations to oscillations in $\Lambda(t)$ will always go to zero eventually. When this happens, the dynamics of $\Lambda(t)$ become deterministic. Conversely, if $\epsilon\geq 1$, then all oscillations will eventually fall below the background fluctuation level. the trajectory of $\Lambda(t)$ is dominated by noise and local oscillations caused by spontaneous cell mass synchronicity. In Appendix E we show how in the limit of vanishing growth rate correlations, the condition $\epsilon<1$ is equivalent to Eq. (1).
In Fig. 1 we show examples of the trajectories of $\Lambda(t)$ for $\epsilon=0.5$ and $\epsilon=1.3$, as well as the log of the deviation of $\Lambda(t)$ from its asymptotic value $\Lambda_{\infty}$. The oscillation fit and envelope are obtained by taking the theoretical transient oscillatory behavior and matching the amplitude and phase to the cell mass distribution at the end of the simulation. The fluctuation envelope scales with $\sqrt{N(t)}$ (see Appx. C for precise definitions and derivations of the oscillation and fluctuation envelopes). In log deviation space, the exponential decay rates of transient effects appear as slopes. Which transient effect dominates in the large time limit only depends on the theoretical ratio of the slopes set by $\epsilon$. One can think of the point at which the oscillations clearly separate from the background noise as the crossover between the two lines. At the transition, as $\epsilon$ goes to $1$, this separation time diverges. In Sec. IV we will show that this dynamic transition at $\epsilon=1$ has properties that are analogous to equilibrium phase transitions.